# Synthesizing High-Quality Programming Tasks with LLM-based Expert and Student Agents

### v1

@@ Abstract
Generative AI is transforming computing education by enabling the automatic generation of personalized content and feedback. We investigate its capabilities in providing high-quality programming tasks to students. Despite promising advancements in task generation, a quality gap remains between AI-generated and expert-created tasks. The AI-generated tasks may not align with target programming concepts, could be incomprehensible for students to solve, or may contain critical issues such as incorrect tests. Existing works often require interventions from human teachers for validation. We address these challenges by introducing PyTaskSyn, a novel synthesis technique that first generates a programming task and then decides whether it meets certain quality criteria to be given to students. The key idea is to break this process into multiple stages performed by expert and student agents simulated using both strong and weaker generative models. Through extensive evaluation, we show that PyTaskSyn significantly improves task quality compared to baseline techniques and showcases the importance of each specialized agent type in our validation pipeline. Additionally, we conducted user studies using our publicly available web application and show that PyTaskSyn can deliver high-quality programming tasks comparable to expert-designed ones while reducing workload and costs, and being more engaging than programming tasks that are available in online resources.
@@ 1 Introduction
Generative AI is transforming learning and teaching in computing education [1, 2, 3]. Advanced generative models such as OpenAI’s GPT-4o [4] and GitHub Copilot [5] are reshaping both student and teacher experiences. For students, these models can provide personalized educational content [6], provide detailed feedback on their work [7, 8], and serve as pair programmers [9, 10]. For teachers, these models can assist in analyzing and grading student answers [11] and curriculum development [12]. A particularly promising application is their ability to generate tailored educational materials, especially in creating diverse programming exercises that target specific programming concepts.
Recent works have investigated the use of generative models for generating novel and engaging programming exercises related to a specific theme and targeting specific programming concepts [13, 14, 15, 16]. While these initial efforts show promise, AI-generated tasks still fall short of human expert quality due to several issues [14, 11, 16]. For example, the generated programming task can contain incorrect test cases generated as part of the task or it can contain a task description that is not comprehensible [14, 16]. Without an automatic validation mechanism to check these aspects of the task, human interventions would be required to validate the task’s quality before they are assigned to students [17, 14]. While one might consider using generative models for task validation, research has shown that they struggle with self-correction [18]. These limitations in single-agent validation motivated our multi-agent based approach.
To address these challenges, we introduce our technique, PyTaskSyn, capable of generating contextualized programming tasks and then deciding whether they meet certain quality criteria. Figure 1 depicts the overview of the pipeline implemented in PyTaskSyn. First, given a context as input, PyTaskSyn asks a simulated agent to generate a programming task composed of a description and a test suite. The generated task then goes through the second stage in the pipeline (i.e., validation), handled by multiple simulated agents with unique roles using strong (GPT-4o [4]) and weaker (GPT-4o-mini [19]) models. This validation stage is designed to provide a quality assurance mechanism and decides whether the generated task can be provided to a student or not. Our multi-agent approach builds upon research that has shown the benefit of collaborative agents [20, 21, 22] and used generative models to simulate humans in various roles [23, 24, 25, 8]. We demonstrate the efficacy of our technique through extensive evaluation using data collected from prior works [13, 17, 15, 16].
@@
    We highlight the effectiveness of decomposing the programming task synthesis into multiple stages. We introduce PyTaskSyn, a novel technique that leverages generative models to simulate expert, tutor, and student agents, each responsible for a specific stage (Section 4).
@@ 2 Related Work
Programming task generation. With the advancement of generative models, recent studies have demonstrated the potential of leveraging them to create natural language programming tasks [13, 17, 14, 15, 11, 16]. Initial research focused on the generation of programming exercises, including sample solutions and test cases [16], while subsequent research efforts have aimed at synthesizing more diverse and contextualized programming exercises [13, 17, 15], as well as exercises suitable for non-English contexts [14]. A parallel line of work focused on generating programming tasks to target specific types of bugs [11], or visual programming tasks [26, 27, 28]. However, many of these works highlighted the significant gap between AI-generated tasks and human-crafted tasks [14, 27, 11, 16]. Our work addresses this gap by automating the validation of programming tasks’ quality.
Improving quality of programming content. Recent works focused on enhancing the quality of programming content. For example, generative models were used to generate predicates for testing generated programs [29], or to act as judges for evaluation [30]. As an enhancement, they can leverage their own evaluations to iteratively improve the quality of their responses [31, 32]. However, prior work has shown that validating the feedback generated by one generative agent using another generative agent has proved to be more effective than using the same agent [25, 8]. Similarly, in our work, we break the task synthesis process into smaller processes and assign them to different agents.
Generative models as simulated agents. A recent line of research involves agents assuming different roles and collaborating to achieve a goal. For example, in AutoGen [22], multiple agents can converse, use tools, and incorporate human input. Similar studies show that simulated agents interacting with each other outperform a single agent in reasoning and planning [20, 21]. Several studies have explored LLM-empowered agents for simulating classroom interactions [33], facilitating tutor training [34], and implementing learning-by-teaching environments [35, 36, 37, 38]. Our work builds upon these successes by simulating experts, tutors, and students while synthesizing tasks to ensure their quality.
@@ 3 Problem Setup
Programming concepts: Dictionaries, Classes &\&& Objects, Strings, Arithmetic Operators
@@
Figure 2: Example of a contextualized programming task. The task description provides comprehensible information for solving the task. However, the task contains an incorrect test and fails to cover the “Dictionaries” concept. PyTaskSyn validates the quality of this task and abstains from outputting it to students.
@@ 3.2 Quality of Contextualized Programming Tasks
Ensuring the quality of contextualized programming tasks is crucial before giving them to students. These tasks must be relevant to the desired theme and programming concepts while being correct and comprehensible for students to solve. We propose a systematic evaluation process for a human expert to assess the quality of task $\mathcal{T}$ created for context $\psi$. The process begins with the expert formulating a solution for $\mathcal{T}$, which is a natural approach to gain a thorough understanding of the task. If the expert cannot formulate a solution, this indicates a fundamental issue with the task, marking it as low-quality. Upon successfully formulating a solution code $\mathcal{C}^{*}$, the expert evaluates the correctness of $\mathcal{T}$ by verifying whether $\mathcal{T}_{\mathrm{tests}}$ correctly validates $\mathcal{C}^{*}$, covering all base and corner cases handled by $\mathcal{C}^{*}$. Next, they verify whether task $\mathcal{T}$ is relevant to $\psi_{\mathrm{theme}}$ and whether all the programming concepts in $\psi_{\mathrm{concepts}}$ are required when formulating $\mathcal{C}^{*}$. Then, they assess whether the task description $\mathcal{T}_{\mathrm{desc}}$ provides sufficient information for students to write solutions. We define Q-Overall as the overall quality metric, where an expert assigns a final numerical score at the end of the evaluation process. To reduce potential biases, we involve two experts and aggregate their scores for our experimental evaluation in Section 5.
@@ 3.3 Technique Evaluation Metrics
To guide our technique development and evaluate its effectiveness, we use two performance metrics: (i) Coverage, measuring the percentage of times our technique provides a programming task to the student, and (ii) Precision, measuring the percentage of times the programming task provided to the student is of high quality. Our objective is to develop a technique with high precision to ensure well-designed tasks and high coverage rate to provide tasks across diverse contexts.
@@ 4 Our Technique PyTaskSyn
a-b). If the task does not pass the validation stage, our technique retries up to $N$ times, before ultimately deciding whether or not to assign a task to the student.
4.1 Motivation and Overview
Common validation techniques use a single generative agent to generate both a task and its solution code for consistency checking [16] or to act as a judge for evaluating the output task [30]. However, our initial experiments revealed task quality issues when relying on a single agent (cf. Section 5.4). Previous research show that breaking a complex process into sub-processes and combining the capabilities of multiple agents in a modular manner significantly improves performance [20, 22]. Building on this insight, we implement a multi-agent technique where each agent takes a unique role. Figure 3 presents an overview of our technique consisting of a generation stage and a validation stage. The validation aims to improve precision, but may result in less coverage. To tackle the coverage drop, we use an outer loop. Concretely, we run the synthesis process until one task passes validation or we reach a maximum number of trials, $N$. If no task passes after $N$ attempts, we do not provide any task to the student.
@@ 4.3 Stage 2 - Validation
Stage 2a - Validation by SimTutor. In this stage, our goal is to evaluate the test suite and contextual relevance of the task generated in Stage 1. To this end, we implement a SimTutor agent to simulate a human tutor for validating the task. Figure 4 (Stage 2a) shows the prompt we use for this SimTutor agent. It is assigned the role of a tutor through a system prompt, followed by a user prompt containing an input context $(\psi_{\text{theme}},\psi_{\text{concepts}})$, and a generated task $\mathcal{T}$. First, we instruct the SimTutor agent to solve the task by writing a solution code $\mathcal{C}^{*}$. This approach follows the Chain-of-Thought prompting strategy [39], which mimics the evaluating process of a human expert in Section 3.2. Solution $\mathcal{C}^{*}$ is then used to validate $\mathcal{T}_{\mathrm{tests}}$, since even state-of-the-art generative models such as GPT-4o can produce incorrect test cases. Specifically, we verify whether all test cases pass when executing $\mathcal{C}^{*}$ and whether every line of $\mathcal{C}^{*}$ is covered by $\mathcal{T}_{\mathrm{tests}}$. Then, the agent assesses the relevance of the task to the given context $\psi$. It assigns a score of 1 if task $\mathcal{T}$ effectively integrates the given theme and programming concepts, and a score of 0 otherwise. We note that if the agent attempts to cheat by writing a code $\mathcal{C}^{*}$ that directly uses input-output pairs from $\mathcal{T}_{\mathrm{tests}}$, the context relevance score will be 0 as it fails to use the required programming concepts. We use the GPT-4o model with its default temperature of 1.0 for our SimTutor agent.
@@ 5.1 Context Selection
We evaluate the effectiveness of our technique across varied themes and programming concepts collected from prior works [13, 17, 15, 16]. We select $5$ diverse themes and uniformly sample $5$ sets of $3$ to $5$ core Python programming concepts for each theme, resulting in $25$ contexts in total. Examples of themes and programming concepts are shown in Figure 7.
@@ 5.2 Evaluation Procedure
For each sampled context $\psi$, we generate $N=10$ programming tasks, resulting in a pool of $250$ tasks ($25$ contexts $\times$ $10$ tasks). This is done prior to applying different validation techniques in Section 5.3. We involve two human annotators who are experts in Computer Science and Python programming to evaluate the tasks using the Q-Overall metric introduced in Section 3.2 and assign binary scores(1-High quality/0-Low quality) to each task. [^1] Additionally, our expert evaluation shows that 98% of the 250 generated tasks do not contain programming concepts more advanced than those in the input context. To better understand the reasoning behind evaluations, we ask the two annotators to answer three additional Yes (1)/No (0) questions: (i) Is the test suite correct and sufficiently covering relevant cases?; (ii) Does the task accurately reflect the input context?; (iii) Is the task description comprehensible? We obtain a Cohen’s Kappa agreement [40] of $0.8$ for the Q-Overall metric and at least $0.7$ for each additional question, indicating substantial agreement between the annotators. We aggregate the two annotators’ scores to obtain average quality scores for each task. Tasks passed validation of each technique in Section 5.3 are used for computing precision and coverage.
@@ 5.3 Techniques Evaluated
Baselines. First, Base neither applies a consistency check during the generation stage nor uses any validation mechanisms in the validation stage. Second, GenConsistency incorporates only the consistency check during the generation stage [16]. Third, LLMJudge leverages an LLM as a judge [30] to validate a task in the validation stage. We prompt it with input context $\psi$, a task $\mathcal{T}$, then instruct it to assess the test suite, contextual relevance, and task comprehensibility. LLMJudge assigns a binary Q-Overall score (1/0) for the task’s overall quality.
@@
Figure 5: Technique comparison. $\textsc{PyTaskSyn}_{\tau}$ improves task quality substantially over baselines GenConsistency and LLMJudge, getting closer to the performance of $\textsc{Oracle}_{p}$. We vary $\tau\in\{0\%,50\%,100\%\}$ to showcase the tunable precision-coverage trade-off of $\textsc{PyTaskSyn}_{\tau}$.
@@ 6 User Studies Using Our Web Application
We created a publicly available web application where users can request tasks from PyTaskSyn for their chosen themes and programming concepts, write code and debug using our integrated programming environment. [^2] Link to our web application is anonymized for review. We conducted two user studies via our web application to compare tasks synthesized by PyTaskSyn against tasks from other sources (Section 6.1), and to evaluate the real-world performance of our web application (Section 6.2).
@@ 6.1 Comparison with Expert-created Tasks and Online Resources
Solving
time
Creation
time
Expert $0.95$ $0.95$ $0.95$ $0.42$ $0.72$ 184848484 $10.82$ $25.20$
[…]
Figure 8: Comparison of tasks from expert, online resources, and PyTaskSyn. We report averaged participant ratings on task aspects and other statistics. Tasks from PyTaskSyn achieved quality comparable to expert-designed ones while requiring substantially less creation time. Tasks from online resources lacked thematic integration and were rated as less engaging and not challenging.
Setup. We compared the quality of tasks synthesized by PyTaskSyn against tasks created by experts and those from online resources. From 5 themes in Section 5.1, we re-sampled $3$ to $5$ programming concepts to create 5 new contexts. For each context, we collected $3$ tasks from: (1) our web application, (2) an expert, and (3) online resources (specifically geeksforgeeks.org). [^3] We made our best effort to find tasks on geeksforgeeks.org that covered programming concepts in each context. The presence of specific themes generally cannot be found. We recruited $10$ participants, including tutors and graduate students who volunteered for our study. They are non-native English speakers, with an average age of $28.3$. Among them, $5$ held Master’s degrees and $5$ held Bachelor’s degrees in STEM fields, with an average of $6.8$ years of Python programming experience. Each participant received $2$ contexts with their corresponding $6$ tasks, in a blind condition where task sources were not disclosed. After either completing each task or reaching the $20$-minute time limit, they provided feedback using multi-level Likert scales on: theme relevance (1-Yes/0.5-Partially/0-No), programming concepts relevance (1-Yes/0.5-Partially/0-No), task comprehensibility (1-Yes/0-No), difficulty (1-Hard /0.5-Medium/0-Easy), and interestingness (1-Interesting/0.5-Okay/0-Boring).

### v2

@@ Abstract
Generative AI is transforming computing education by enabling the automatic generation of personalized content and feedback. We investigate its capabilities in providing high-quality programming tasks to students. Despite promising advancements in task generation, a quality gap remains between AI-generated and expert-created tasks. The AI-generated tasks may not align with target programming concepts, could be incomprehensible to students, or may contain critical issues such as incorrect tests. Existing works often require interventions from human teachers for validation. We address these challenges by introducing PyTaskSyn, a novel synthesis technique that first generates a programming task and then decides whether it meets certain quality criteria to be given to students. The key idea is to break this process into multiple stages performed by expert and student agents simulated using both strong and weaker generative models. Through extensive evaluation, we show that PyTaskSyn significantly improves task quality compared to baseline techniques and showcases the importance of each specialized agent type in our validation pipeline. Additionally, we conducted user studies using our publicly available web application and show that PyTaskSyn can deliver high-quality programming tasks comparable to expert-designed ones while reducing workload and costs, and being more engaging than programming tasks that are available in online resources.
@@ 1 Introduction
Generative AI is transforming learning and teaching in computing education [1, 2]. Advanced generative models such as OpenAI’s GPT-4o [3] and GitHub Copilot [4] are quickly reshaping both student and teacher experiences. For students, these models can provide personalized educational content [5], provide detailed feedback on their work [6, 7], and serve as pair programmers [8]. For teachers, these models can assist in analyzing and grading student answers [9] and curriculum development [10]. A particularly promising application is their ability to generate tailored educational materials, especially in creating diverse programming exercises that target specific concepts.
Recent works have investigated the use of generative models for generating novel and engaging programming exercises related to a specific theme and targeting specific programming concepts [11, 12, 13, 14, 15]. While these initial efforts show promise, AI-generated tasks still fall short of human expert quality due to several issues [12, 9, 14]. For example, the generated programming task can contain incorrect test cases generated as part of the task or it can contain a task description that is not comprehensible [12, 14]. Without an automatic validation mechanism to check these aspects of the task, human interventions would be required to validate the task’s quality before they are assigned to students [16, 12]. While one might consider using generative models for task validation, research has shown that they struggle with self-correction [17]. These limitations in single-agent validation motivated our multi-agent based approach.
To address these challenges, we introduce our technique, PyTaskSyn, capable of generating contextualized programming tasks and then deciding whether they meet certain quality criteria. Figure 1 depicts the overview of the pipeline implemented in PyTaskSyn. First, given a context as input, PyTaskSyn asks a simulated agent to generate a programming task composed of a description and a test suite. The generated task then goes through the second stage in the pipeline (i.e., validation), handled by multiple simulated agents with unique roles using strong (GPT-4o [3]) and weaker (GPT-4o-mini [18]) models. This validation stage is designed to provide a quality assurance mechanism and decides whether the generated task can be provided to a student or not. Our multi-agent approach builds upon research that has shown the benefit of collaborative agents [19, 20, 21] and used generative models to simulate humans in various roles [22, 23, 24, 7, 25]. We demonstrate the efficacy of our technique through extensive evaluation using data collected from prior works [11, 16, 13, 14].
@@
    We highlight the effectiveness of decomposing the programming task synthesis into multiple stages. We introduce PyTaskSyn, a novel technique that leverages generative models to simulate expert, tutor, and student agents, each responsible for specific stages in the synthesis process. (Section 4). [^1] https://github.com/machine-teaching-group/aied2025-pytasksyn
@@
2 Related Work
Programming task generation. With the advancement of generative models, recent studies have demonstrated the potential of using them to create natural language programming tasks [11, 16, 12, 13, 9, 14]. Initial research focused on the generation of programming exercises, including sample solutions and test cases [14], while subsequent research efforts have aimed at synthesizing more diverse and contextualized programming exercises [11, 16, 13], as well as exercises suitable for non-English contexts [12]. A parallel line of work focused on generating programming tasks to target specific types of bugs [9], or visual programming tasks [26, 27, 28]. However, many of these works highlighted the significant gap between AI-generated tasks and human-crafted tasks [12, 27, 9, 14]. Our work addresses this gap by automating the validation of programming tasks’ quality.
Improving quality of programming content. Recent works have focused on enhancing the quality of programming content. For example, generative models were used to generate predicates for testing a generated program [29], or to act as judges for evaluation [30]. As an enhancement, generative models can leverage their own evaluations to iteratively improve the quality of their responses [31, 32]. However, prior work has shown that validating the feedback generated by one generative agent using another generative agent has proved to be more effective than using the same agent [24, 7]. Similarly, in our work, we break the task synthesis process into smaller processes and assign them to different agents.
Generative models as simulated agents. A recent line of research involves agents assuming different roles and collaborating to achieve a goal. For example, in AutoGen [21], multiple agents can converse, use tools, and incorporate human input. Similar studies show that simulated agents interacting with each other outperform a single agent in reasoning and planning [19, 20]. Several studies have explored LLM-empowered agents for simulating classroom interactions [33], facilitating tutor training [34], and implementing learning-by-teaching environments [35, 36, 37]. Our work builds upon these successes by simulating experts, tutors, and students while synthesizing tasks to ensure their quality.
@@ 3 Problem Setup
Programming concepts: Dictionaries, Classes &\& Objects, Strings, Arithmetic Operators
@@
Figure 2: Example of a contextualized programming task. This task has an incorrect test and fails to cover the “Dictionaries” concept. Our technique PyTaskSyn validates the quality of this task and abstains from outputting it to students.
@@ 3.2 Quality of Contextualized Programming Tasks
Ensuring the quality of a contextualized programming task is crucial before giving it to students. It must be relevant to the desired theme and programming concepts while being correct and comprehensible for students to solve. We propose a systematic evaluation process for a human expert to assess the quality of task $\mathcal{T}$ created for context $\psi$. The process begins with the expert formulating a solution for $\mathcal{T}$, which is a natural approach to gain a thorough understanding of the task. If the expert cannot formulate a solution, this indicates a fundamental issue with the task, marking it as low-quality. Upon successfully formulating a solution code $\mathcal{C}^{*}$, the expert evaluates the correctness of $\mathcal{T}$ by verifying whether $\mathcal{T}_{\mathrm{tests}}$ correctly validates $\mathcal{C}^{*}$, covering all base and corner cases handled by $\mathcal{C}^{*}$. Next, they verify whether task $\mathcal{T}$ is relevant to $\psi_{\mathrm{theme}}$ and whether all the programming concepts in $\psi_{\mathrm{concepts}}$ are required when formulating $\mathcal{C}^{*}$. Then, they assess whether the task description $\mathcal{T}_{\mathrm{desc}}$ provides sufficient information for students to write solutions. We define Q-Overall as the overall quality metric, where an expert assigns a final numerical score at the end of the evaluation process.
@@ 3.3 Technique Evaluation Metrics
To guide our technique development and evaluate its effectiveness, we use two performance metrics: (i) Coverage, measuring the percentage of times a technique provides a programming task to the student, and (ii) Precision, measuring the percentage of times the programming task provided to the student is of high quality. Our objective is to develop a technique with high precision to ensure well-designed tasks and high coverage rate to provide tasks across diverse contexts.
@@ 4 Our Technique PyTaskSyn
4.1 Motivation and Overview
Common validation techniques use a single generative agent to generate both a task and its solution code for consistency checking [14] or to act as a judge for evaluating the output task [30]. However, our initial experiments revealed task quality issues when relying on a single agent (cf. Section 5.4). Previous research show that breaking a complex process into sub-processes and combining the capabilities of multiple agents in a modular manner significantly improves performance [19, 21]. Building on this insight, we implement a multi-agent technique where each agent takes a unique role. Figure 3 presents an overview of our technique consisting of a generation stage and a validation stage. The validation aims to improve precision, but may result in less coverage. To tackle coverage drop, we use an outer loop that repeats synthesis until a task passes validation or a maximum of $N$ trials is reached; if none pass, no task is given to the student.
@@ 4.1 Motivation and Overview
a-b). If the task does not pass the validation stage, our technique retries up to NN times, before ultimately deciding whether or not to assign a task to the student.
@@ 4.3 Stage 2 - Validation
Stage 2a - Validation by SimTutor. In this stage, our goal is to evaluate the test suite and contextual relevance of the task generated in Stage 1. To this end, we implement a SimTutor agent to simulate a human tutor for validating the task. Figure 4 (Stage 2a) shows the prompt we use for this SimTutor agent. It is assigned the role of a tutor through a system prompt, followed by a user prompt containing an input context $(\psi_{\text{theme}},\psi_{\text{concepts}})$, and a generated task $\mathcal{T}$. First, we instruct the SimTutor agent to solve the task by writing a solution code $\mathcal{C}^{*}$. This approach follows the Chain-of-Thought prompting strategy [38], which mimics the evaluating process of a human expert in Section 3.2. Solution $\mathcal{C}^{*}$ is then used to validate $\mathcal{T}_{\mathrm{tests}}$. Specifically, we verify whether all test cases pass when executing $\mathcal{C}^{*}$ and whether every line of $\mathcal{C}^{*}$ is covered by $\mathcal{T}_{\mathrm{tests}}$. Then, the agent assesses the relevance of the task to the given context. It assigns a score of 1 if task $\mathcal{T}$ effectively integrates the given theme and programming concepts, and a score of 0 otherwise. We note that if the agent attempts to cheat by writing a code $\mathcal{C}^{*}$ that directly uses input-output pairs from $\mathcal{T}_{\mathrm{tests}}$, the context relevance score will be 0 as it fails to use the required programming concepts. We use the GPT-4o model with its default temperature of 1.0 for our SimTutor agent.
@@ 5.1 Context Selection
We evaluate the effectiveness of our technique across varied themes and programming concepts collected from prior works [11, 16, 13, 14]. We select $5$ diverse themes and uniformly sample $5$ sets of $3$ to $5$ core Python programming concepts for each theme, resulting in $25$ contexts in total (cf. Figure 7 for examples).
@@ 5.2 Evaluation Procedure
For each sampled context $\psi$, we generate $N=10$ programming tasks, resulting in a pool of $250$ tasks ($25$ contexts $\times$ $10$ tasks). This is done prior to applying different validation techniques in Section 5.3. Two authors, with expertise in Python programming, evaluated the tasks using the Q-Overall metric introduced in Section 3.2 and assigned binary scores (1-High quality/0-Low quality) to each task. [^2] Additionally, our expert evaluation shows that 98% of the 250 generated tasks do not contain programming concepts more advanced than those in the input context. To better understand the reasoning behind evaluations, we ask the two annotators to answer three additional Yes(1)/No(0) questions: (i) Is the test suite correct and sufficiently covering relevant cases?; (ii) Does the task accurately reflect the input context?; (iii) Is the task description comprehensible? We obtain a Cohen’s Kappa agreement [39] of $0.8$ for Q-Overall metric and at least $0.7$ for each additional question, indicating substantial agreement between the annotators. We aggregate the two annotators’ scores to obtain average quality scores for each task. Tasks passed validation of each technique in Section 5.3 are used for computing precision and coverage.
@@ 5.3 Techniques Evaluated
Baselines. First, Base neither applies a consistency check during the generation stage nor uses any validation mechanisms in the validation stage. Second, GenConsistency incorporates only a consistency check during the generation stage [14]. Third, LLMJudge leverages an LLM as a judge [30] to validate a task in the validation stage. We prompt it with input context $\psi$, a task $\mathcal{T}$, then instruct it to assess the test suite, contextual relevance, and task comprehensibility. It assigns a binary Q-Overall score (1/0) for the task’s overall quality.
@@
Figure 5: Technique comparison. $\textsc{PyTaskSyn}_{\tau}$ improves task quality substantially over both baselines GenConsistency and LLMJudge, getting closer to the performance of $\textsc{Oracle}_{p}$. We vary threshold $\tau\in\{0\%,50\%,100\%\}$ to showcase the tunable precision-coverage trade-off of $\textsc{PyTaskSyn}_{\tau}$.
@@ 6 User Studies Using Our Web Application
We created a public web application (https://pytasksyn.netlify.app) where users can request tasks from PyTaskSyn for their chosen themes and programming concepts, write code and debug using our integrated programming environment. We conducted the following two user studies via this web application.
@@ 6.1 Comparison with Expert-created Tasks and Online Resources
Setup. We compared tasks from PyTaskSyn against tasks created by experts and those from online resources. From 5 themes in Section 5.1, we re-sampled $3$ to $5$ programming concepts to create 5 new contexts. For each, we collected $3$ tasks from: (1) an expert, (2) online resources (specifically geeksforgeeks.org), and (3) our web application. [^3] We made our best effort to find tasks on geeksforgeeks.org that covered programming concepts in each context. The presence of specific themes generally cannot be found. We recruited $10$ volunteer participants, including tutors and graduate students. They are non-native English speakers, with an average age of $28.3$. Among them, $5$ held Master’s degrees and $5$ held Bachelor’s degrees in STEM fields, with an average of $6.8$ years of Python experience. We assign each participant $2$ random contexts with their corresponding $6$ tasks (task sources are undisclosed). After completing each task or reaching the $20$-minute limit, they provided feedback using multi-level Likert scales on: theme relevance (1-Yes/0.5-Partial/0-No), programming concepts relevance (1-Yes/0.5-Partial/0-No), task comprehensibility (1-Yes/0-No), difficulty (1-Hard/0.5-Medium/0-Easy), and interestingness (1-Interesting/0.5-Okay/0-Boring).
@@
Solving time
Creation time
Expert $0.95$ $0.95$ $0.95$ $0.42$ $0.72$ 18484 $10.82$ $25.20$
Online resources $0.15$ $0.82$ $0.85$ $0.15$ $0.50$ $100$ $\phantom{1}3.26$ 01
PyTaskSyn $0.98$ $0.92$ $0.90$ $0.20$ $0.65$ $100$ $\phantom{1}6.66$ 11.291.29
Figure 8: Comparison of tasks from expert, online resources, and PyTaskSyn. We report averaged participant ratings on various task aspects and other statistics. Tasks from PyTaskSyn achieved quality comparable to expert-designed ones while requiring substantially less creation time.
Statistic Participant ID Average
P1 P2 P3 P4 P5 P6 P7 P8 P9 P10
No. Requests 66 55 66 66 55 55 66 55 55 55 $5.40$
No. Synthesized Tasks 55 55 55 55 55 55 55 55 55 55 $5.00$
No. Solved Tasks 55 55 44 44 44 55 55 55 33 55 $4.50$
Figure 9: Statistics of each participant when they were instructed to request and solve tasks synthesized by PyTaskSyn on our web application. Our system successfully synthesized tasks for 92.6% of participant requests.